OpenStudy (anonymous):

Write the exponential equation in logarithmic form. 3x = 243 1.log3243 = x 2.log243^x = 3 3.logx^3 = 243 4.log3^x = 243

OpenStudy (anonymous):

if you dont get the answer just give me your answer. thx

OpenStudy (anonymous):

Is the original equation \(3^x = 243\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log base3 of 243 = x

OpenStudy (anonymous):

Ok, so remember the definition of the logarithm: \[b^k = a \iff log_b (a) = k\] Look at your problem, and the part of the definition on the left side. From that general description, what would be the b, the k, and the a in your problem?

OpenStudy (anonymous):

this has always been hardest for me...

OpenStudy (anonymous):

\[3^x = 243\] \[b^k = a\] If I tell you that those two equations are equivalent, what is the 'b' in the first equation?

OpenStudy (anonymous):

u might find it easier to convert this type of espression by rememberibg 10^ 2 = 100 10 is the base, 2 is the log so log to the base 10 of 100 = 2 the log is the power

OpenStudy (anonymous):

sometimes it is easier to remember the relation in numbers than in letters

OpenStudy (anonymous):

I find the opposite to be true. Numbers just confuse what I'm doing since those typically get plugged into the calculator and lose all meaning. ;p

OpenStudy (anonymous):

yes - it depends on the person- my teaching experience has taught me that

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

if it is 3^x = 243 then \[\log_{3 } 243 = x\]

OpenStudy (anonymous):

I think (1) is true. but it is written a little untidy.